prove that the function f given by


f(x)={ 2,if x is irrational


-2,if x is rational


is discontinuous, for all x€R,using the sequential definition of continuity

1
Expert's answer
2022-01-14T04:18:05-0500

ANSWER.

By the definition of continuity from condition "\lim _{ n\rightarrow \infty }{ { x }_{ n } } =a" it follows "\lim _{ n\rightarrow \infty }{ { f(x }_{ n })=f(a) } \\ \\" . Therefore , if "\lim _{ n\rightarrow \infty }{ { x }_{ n } } =\lim _{ n\rightarrow \infty }{ { y }_{ n } } =a" then "\lim _{ n\rightarrow \infty }{ { f(x }_{ n })=\lim _{ n\rightarrow \infty }{ { f(\ y }_{ n })\ \ } }"

Let "a" be a real number and "n" be a natural number. Between the numbers "a-1\/n , a+1\/n" we choose a rational number "x_n" and irrational number "y_n" . Since "a-\frac { 1 }{ n } <{ x }_{ n }<a+\frac { 1 }{ n } \quad ,a-\frac { 1 }{ n } <{ y }_{ n }<a+\frac { 1 }{ n }" , and "\lim _{ n\rightarrow \infty }{ \left( a-\frac { 1 }{ n } \right) = } \lim _{ n\rightarrow \infty }{ \left( a+\frac { 1 }{ n } \right) =a }", then "\lim _{ n\rightarrow \infty }{ { x }_{ n }= } \lim _{ n\rightarrow \infty }{ { y }_{ n }=a }" . For sequences "\left( { x }_{ n } \right) \ in\ Q,\ \left( { y }_{ n } \right) \ in\ R\setminus Q" we have "f\left( { x }_{ n } \right) =-2,\ f\left( { y }_{ n } \right) = 2" for all "n\in N" . So

"\lim _{ n\rightarrow \infty }{ { f(x }_{ n })=-2\neq } 2=\ \lim _{ n\rightarrow \infty }{ { f(\ y }_{ n })\ \ } ."

Hence, "f" is not continuous at "a" , and as "a" was arbitrary , that "f" is not continuous at any "a\in R."